Reports 1-1 of 1 Clear search Modify search
fabian.arellano - 16:22 Wednesday 27 October 2021 (18710) Print this report
Comment to BS heigth measurement (17338)

With Hirata-san.

Summary: the BS optic is 2.8 mm below the average height of the flanges when F0 and BF are close to zero and F1 is alledgely close to zero.

See pictures in the album BS Remedying work after O3.

  • Temperature: T_ondotori = 23.6  °C, T_PT100 = 23.6  °C.
  • We set the F0 anbd BF GAS fiters at zero with coil magnet actuation:
    • F0: 14,000 cnt,
    • BF: 4000 cnt,
  • F1 LVDT is not working well (entry 18191), so we checked the amount of actuation we used the last time we measured the height, when the LVDT was working, and used the same value (entry 17338 from the 30th of June, 2021). This is fine because we have the same temperature today (23.6 °C) and the fishing rod has not been used:
    • F1 actuation: -10,000 cnt (per Data Viewer).
    • F1 LVDT: 787 um (nominal is zero but we have a faulty LVDT primary coil).
  • As a reference we used Tomaru-san's marking of the average of the heights of the four flanges of the chamber.
  • The measured height to be approximately 2.8 mm below the reference.
Search Help